OpenStudy (anonymous):

A new athletic shoe store had expenses of $40,000 for designing and building the shelves and counters and $120,000 for the first year’s shoe inventory. So far this year, the shoe sales are $70,000. What do the sales have to be for the rest of the year for the store to break even? A. $40,000 B. $70,000 C. $80,000 D. $90,000

OpenStudy (anonymous):

@pierodog

OpenStudy (anonymous):

D, again some more simple subtraction but glad to help

OpenStudy (anonymous):

D. 12000 + 40000 = 160000 - 70000 = 90000